a)     Determine the Laplace transforms of the following functions


(i) f(t) = t3-2t

(ii) f(t) = sin3t-e2t

(iii) f(t) = e2t sinh4t



b) Find the Laplace transform of the following

(i) f(t) = 3e-4t- 5e4t

(ii) f(t) = t sin 3t + cos 4t


find the inverse transform of

(i) F (s) = 2/3-1/2s-3

(ii) F (s) = 5s-8/s(s-4), using partial fractions.


1
Expert's answer
2021-12-16T11:21:46-0500

a)     

(i) "f(t) = t^3-2t"


"F(s)=L\{t^3-2t\}"

"=\dfrac{6}{s^4}-\dfrac{2}{s^2}"

ii) "f(t) = \sin(3t)-e^{2t}"


"F(s)=L\{\sin(3t)-e^{2t}\}"

"=\dfrac{3}{s^2+9}-\dfrac{1}{s-2}"

iii) "f(t) = e^{2t}\sinh(4t)"


"F(s)=L\{e^{2t}\sinh(4t)\}"

"=\dfrac{4}{(s-2)^2-16}"

b)

(i) "f(t) = 3e^{-4t}- 5e^{4t}"


"F(s)=L\{3e^{-4t}- 5e^{4t}\}"

"=\dfrac{3}{s+4}-\dfrac{5}{s-4}"

(ii) "f(t) = t\sin(3t)+\cos(4t)"


"F(s)=L\{ t\sin(3t)+\cos(4t)\}"

"=\dfrac{6s}{(s^2+9)^2}-\dfrac{s}{s^2+16}"

c)

(i) "F (s) =\dfrac{2}{3}-\dfrac{1}{2s-3}"


"f(t)=L^{-1}\{\dfrac{2}{3}-\dfrac{1}{2s-3}\}"

"=\dfrac{2}{3}\delta(t)-\dfrac{1}{2}e^{3t/2}"

(ii) "F (s) =\dfrac{5s-8}{s(s-4)}"


"\dfrac{5s-8}{s(s-4)}=\dfrac{A}{s}+\dfrac{B}{s-4}=\dfrac{A(s-4)+Bs}{s(s-4)}"

"s=0:-4A=-8=>A=2"

"s=0:4B=12=>B=3"

"f(t)=L^{-1}\{\dfrac{5s-8}{s(s-4)}\}=L^{-1}\{\dfrac{2}{s}\}+L^{-1}\{\dfrac{3}{s-4}\}"

"=2+3e^{4t}"